A 110.0 pound patient is to receive Heparin by IV with an infusion rate of 18.0 Units/kg/hour. Heparin is supplied at 25,000 Units in 250.0 mL. How many milliliters per hour should the IV pump be set for this infusion?
Answer #1

ANSWER.

* Calculation of ml/hr.

        ml/hr = Volume(ml) / Time(hr).

     Calculation of volume

        Volume(ml) = Desire dose(D) /Dose in hand(H) * Quantity(Q)

      Here

           Ordered dose = 18 unit/kg/hr.

           Weight of the patient = 110 pound.

     Convert pound to kg

                   1 pound = 0.454 kg.

        So

            Weight of the patient = 110 * 0.454

                                              = 49.9 kg.

        Desired dose (D) for this patient = 18 * 49.9

                                                             = 898 unit/hr.

                  Dose in hand (H) = 25000 units.

                      Quantity (Q) = 250 ml

       As per above formula

            Volume(ml) of drug required = D/H * Q

                                                           = 898/25000 * 250

                                                           = 0.036 * 250

                                                           = 9 ml.